Find the dimensions of a rectangular region with the maximum area that can be enclosed by 60 feet of fencing.
Answer by Cintchr(481)

p = 2(l+w) = 60 ..... A = l * w
Possible dimensions are ....
Perimiter ....... Area
2(1+29) ......... 1 * 29 = 29
2(2+28) ......... 2 * 28 = 56
2(3+27) ......... 3 * 27 = 81
etc ....
notice that the area goes up as you move from a rectangle to a square ...
so if the perimiter is 60 ... then it CAN be a square.
2(15+15) ........ 15 * 15 = 225